# Minimizing the Weighted Number of Tardy Jobs: Data-Driven Heuristic for Single-Machine Scheduling

### v1

@@ 1 Introduction
    Our heuristic consistently surpasses the state-of-the-art approaches (Baptiste et al., 2010; Antonov et al., 2023) in solution quality within the same time limits, achieving a significantly smaller optimality gap and obtaining optimal solutions in 80-100% of cases.
@@ 2.2 Related Data-driven Approaches
Traditional machine learning methods, such as KNN (K-Nearest Neighbors), SVM (Support Vector Machines), decision trees, or shallow neural networks, have been successfully applied to scheduling problems due to their simplicity, broad applicability, and fast performance. These advantages are highlighted in studies on power supply optimization (Saxena et al., 2024; Yang et al., 2022b), with the latter emphasizing the benefits of combining the KNN algorithm and the SVM classifier over more complex neural networks such as LSTM (Long Short-Term Memory). While simpler than deep learning methods, decision trees and perceptrons can effectively address complex scheduling problems. They have found successful applications in multi-mode project scheduling (Portoleau et al., 2024), flexible job-shop and permutation flow-shop problems (Müller et al., 2022; Wang and Tang, 2017), and production line automatic matching (Yang et al., 2022a). In our paper, we also employ traditional machine learning algorithms, as detailed in Sections 5.2 and 6.2.
The application of deep learning algorithms represents a significant area of data-driven methods in scheduling. Based on recent literature, we identify five major trends in how machine learning is employed: (i) learning dispatching rules and heuristics from data (Jun and Lee, 2020; Janssens et al., 2006); (ii) applying reinforcement learning (RL) for scheduling, including policy construction (Monaci et al., 2024; Wu et al., 2024; Yuan et al., 2024; Brammer et al., 2022; Heger and Voss, 2021), metaheuristic control (Alicastro et al., 2021; Zhang et al., 2012), and multi-agent coordination (Liu et al., 2023); (iii) accelerating classical optimization procedures, such as branch-and-price, using predictive models (Koutecká et al., 2024; Václavík et al., 2018; Delgoshaei and Gomes, 2016); (iv) estimating objective values via surrogate models (Bouška et al., 2022); and (v) employing attention-based architectures for task selection and schedule modeling (Du et al., 2024).
@@ 6 Experimental Results
In this section, we first provide an overview of the datasets utilized in our experiments. Subsequently, we describe the conducted experiments, focusing on model selection and comparison with state-of-the-art methods. The proposed algorithm is implemented in Python using the PyTorch library for neural network development. The training labels are obtained on a cluster node (18 Cores/CPU; 2.3GHz; 256 GB RAM), while the tests are performed in Google Colab. The code and data are available at LINK [^1] Link will be available after review.
@@ 6.3 Comparison to the state-of-the-art
500100020003000400050000202040406060808010010010010010010098986262880100100100100100100989869695555Number of jobsInstances solved, %20min60min
@@ 7 Conclusions and Future Work
8 Acknowledgements
This work was funded by the Czech Ministry of Education, Youth and Sports under the ERC CZ project POSTMAN no. LL1902, by the European Union under the project ROBOPROX (reg. no. CZ.02.01.01/00/22_008/0004590) and by the Grant Agency of the Czech Republic under the Project GACR 22-31670S.

### v2

@@ 1 Introduction
    Our heuristic consistently outperforms the state-of-the-art approaches (Baptiste et al., 2010; Antonov et al., 2023) in solution quality within the same time limits, achieving a significantly smaller optimality gap and obtaining optimal solutions in 80-100% of cases.
@@ 2.2 Related Data-driven Approaches
Traditional machine learning methods, such as KNN (K-Nearest Neighbors), SVM (Support Vector Machines), decision trees, or shallow neural networks, have been successfully applied to scheduling problems due to their simplicity, broad applicability, and fast performance. These advantages are illustrated in studies on power supply optimization (Saxena et al., 2024; Yang et al., 2022b), with the latter showing the benefits of combining the KNN algorithm and the SVM classifier over more complex neural networks such as LSTM (Long Short-Term Memory). A similar idea appears in (Módos et al., 2016), where KNN and decision trees are used to estimate unknown time–quality trade-off curves, enabling adaptive control of response times in overloaded systems. In general, traditional ML has been employed in two complementary ways. First, it can be integrated directly into the scheduling process. For instance, decision trees have been applied to predict uncertain activity durations in multi-mode project scheduling (Portoleau et al., 2024); ensembles of trees have been trained to learn effective dispatching rules in flexible job-shop and permutation flow-shop problems (Müller et al., 2022; Wang and Tang, 2017); and shallow neural networks have been used to predict task-to-line assignments in production line matching (Yang et al., 2022a). KNN and SVM classifiers have also demonstrated competitive performance in power supply optimization (Yang et al., 2022b). Second, ML can operate at the meta-level, supporting algorithm selection rather than producing schedules directly. In this line, Chu et al. (2023) develop an SVM-based meta-learning framework that recommends the most suitable metaheuristic for the multi-mode RCPSP, while Uzunoglu et al. (2025) apply a learning-to-rank model to select constructive strategies for serial-batch scheduling, with subsequent improvement through tailored genetic algorithms. While our work also integrates ML into the scheduling process, it differs from earlier studies by using classifiers to predict early/tardy job labels. This application directly addresses the central combinatorial challenge of the $1|\tilde{d}_{i}|\sum w_{i}U_{i}$ problem and enables subsequent refinement through ILP and feasibility checks (see Sections 5.2 and 6.2).
The application of deep learning algorithms represents a significant area of data-driven methods in scheduling. Based on recent literature, we identify five major trends in how machine learning is employed: (i) learning dispatching rules and heuristics from data (Jun and Lee, 2020; Janssens et al., 2006); (ii) applying reinforcement learning (RL) for scheduling, including policy construction (Monaci et al., 2024; Wu et al., 2024; Yuan et al., 2024; Brammer et al., 2022; Heger and Voss, 2021; Wang et al., 2023), metaheuristic control (Alicastro et al., 2021; Zhang et al., 2012), and multi-agent coordination (Liu et al., 2023); (iii) accelerating classical optimization procedures, such as branch-and-price, using predictive models (Koutecká et al., 2024; Václavík et al., 2018; Delgoshaei and Gomes, 2016); (iv) estimating objective values via surrogate models (Bouška et al., 2022); and (v) employing attention-based architectures for task selection and schedule modeling (Du et al., 2024).
@@ 6 Experimental Results
In this section, we first provide an overview of the datasets utilized in our experiments. Subsequently, we describe the conducted experiments, focusing on model selection and comparison with state-of-the-art methods. The proposed algorithm is implemented in Python using the PyTorch library for neural network development. The training labels are obtained on a cluster node (18 Cores/CPU; 2.3GHz; 256 GB RAM), while the tests are performed in Google Colab. The code and data are available at https://doi.org/10.5281/zenodo.17233362 .
@@ 6.3 Comparison to the state-of-the-art
500100020003000400050000505010010010010010010098986262880100100100100100100989869695555Number of jobsInstances solved, %20min60min
@@ 7 Conclusions and Future Work
Acknowledgements
This work was funded by the Czech Ministry of Education, Youth and Sports under the ERC CZ project POSTMAN no. LL1902, by the European Union under the project ROBOPROX (reg. no. CZ.02.01.01/00/22_008/0004590), by the Grant Agency of the Czech Republic under the Project GACR 22-31670S and by the Grant Agency of the Czech Technical University in Prague, grant No. SGS25/144/OHK3/3T/13. Personal note (Nikolai Antonov): To my mom and granny, my beloved Sophia, and my friend Alexander.